# Relative Entropy Methods for Calculating Committors

### v1

@@
Abstract.
Motivated by challenges arising in molecular simulation, we analyze and develop methods of computing reactive trajectories and committor functions for systems described by the overdamped Langevin dynamics. Our main technical advance is a new loss function that measures the accuracy of approximations to the committor function describing a given chemical reaction or other rare transition event. The loss admits a simple interpretation in terms of the distribution of reactive trajectories, and it can be computed in practice to compare the accuracies of different approximations of the committor. We also derive a method of calculating committors by direct minimization of the loss via stochastic gradient descent.
BvK and GE were supported by NSF DMS-2012207
1. Introduction
Motivated by challenges arising in molecular simulation, we analyze and develop methods for computing reactive trajectories and committor functions. To be precise, we consider the stochastic differential equation (2) that characterizes reactive trajectories of the overdamped Langevin dynamics in the transition path theory of Weinan E and Eric Vanden-Eijnden [5]. We assess the effect of replacing the exact committor function $q$ in that equation with an approximation $\tilde{q}$, resulting in approximate reactive trajectories. Our primary theoretical advances are
  1. (1)
    a characterization of those $\tilde{q}$ for which solutions of (2) with $\tilde{q}$ in place of $q$ exist,
  2. (2)
    a characterization of those $\tilde{q}$ for which the distributions of the exact and approximate reactive trajectories are mutually absolutely continuous, and
  3. (3)
    convenient formulas for the change of measure and relative entropy between the exact and approximate distributions for a given $\tilde{q}$.
Surprisingly, the change of measure can be computed exactly without prior knowledge of the exact committor function $q$. Based on this observation, we propose novel computational methods, including
  1. (1)
    a stochastic gradient descent scheme for estimating committor functions that directly minimizes the relative entropy between the exact and approximate distributions of reactive trajectories, and
  2. (2)
    a model assessment procedure for comparing approximate committor functions based on relative entropy differences as in the Bayes and Akaike information criteria.
We limit our study to systems described by the overdamped Langevin dynamics
$\mathrm{d}X_{t}=-\nabla U(X_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}.$ (1)
In models of molecular systems, $X_{t}$ takes values in the space of positions $\mathbb{R}^{3N}$ of the $N$ atomic nuclei comprising the system, $U:\mathbb{R}^{3N}\rightarrow\mathbb{R}$ is a potential energy, and $\varepsilon=k_{B}T$ where $T>0$ is the temperature and $k_{B}$ is Boltzmann’s constant. Each local minimum of $U$ corresponds to a stable state. When the temperature is low, the dynamics is metastable [8]. That is, trajectories spend most of the time vibrating around stable states, undergoing transitions between states only rarely.
Metastability entails severe computational difficulties. One usually wants to estimate statistics of rare transitions, but it may be impractical to compute a trajectory long enough to observe even one. For example, in simulations of proteins, resolving the fastest vibrations requires a time step no larger than $10^{-15}s$, but conformational changes may occur only every $10^{-6}s$ or less frequently. To overcome difficulties related to metastability, enhanced sampling methods have been devised that perturb or condition the dynamics (1) so that transitions occur over short time scales. In various ways, these methods correct for the perturbations to consistently estimate observables of the unperturbed dynamics.
We consider a particular class of enhanced sampling methods based on Transition Path Theory (TPT) [5, 13]. In TPT, one studies transitions between a fixed pair of disjoint sets $A\subset\mathbb{R}^{3N}$ and $B\subset\mathbb{R}^{3N}$, e.g. balls centered at different local minima of $U$ corresponding to different stable states. TPT characterizes reactive trajectories, which are segments of trajectories of (1) observed starting when they last leave $A$ and ending when they first hit $B$; see Figure 1 and Section 2 for a precise definition. The reactive trajectories have the same law as the transition path process
$\mathrm{d}Y_{t}=-\nabla U(X_{t})\,\mathrm{d}t+2\varepsilon\nabla\log q(X_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}$ (2)
where
$q(x)=\mathds{P}[\tau_{B}<\tau_{A}|X_{0}=x]$
@@ 1. Introduction
is the (forward) committor function, i.e. the probability that a trajectory of (1) with $X_{0}=x$ will hit $B$ before $A$. The drift $2\varepsilon\nabla\log q(Y_{t})$ in (2) forces trajectories away from $A$ and towards $B$. As a result, in many cases, transitions between $A$ and $B$ occur much more quickly under (2) than they would under (1); cf. Figure 2.
There has been significant recent interest in sampling reactive trajectories by first computing an approximation $\tilde{q}$ to the committor function and then simulating trajectories of (4) with $\tilde{q}$ in place of $q$; see [17] and numerical methods for computing committor functions such as [6, 7, 3, 9, 12]. Such a procedure can be dramatically more efficient than the direct simulation of long trajectories to observe transitions, but the effect of the approximation of the committor on the law of the reactive trajectories has not been assessed. Our results in Sections 3 and 4 below provide a means of assessing and controlling the quality of the approximation in theory and practice. In Theorem 1, we characterize the set of approximate committor functions $\tilde{q}$ for which the approximate process is absolutely continuous with respect to the exact process, and we derive a convenient formula for the change of measure between the exact and approximate processes.
Surprisingly, the change of measure can be computed explicitly, at least up to a normalizing constant, even when the exact committor function is unknown; cf. equation (11). Considering Girsanov’s theorem, one might expect the change of measure to depend on the difference in drifts $2\varepsilon\nabla\log q(Y_{t})-2\varepsilon\nabla\log\tilde{q}(Y_{t})$. In Section 3, we show that one can totally eliminate the dependence on exact committor $q$. We propose several novel computational methods based on this observation. In Section 4.3, we derive a stochastic gradient descent method for computing committor functions by direct minimization of the relative entropy of the approximate process with respect to the exact process. Competing methods for estimating committors minimize loss functions that do not relate so closely to the law of the approximate transition path process. For example, Ritz methods minimize
$J(\tilde{q}):=\int_{\mathbb{R}^{3N}\setminus(A\cup B)}\lvert\nabla\tilde{q}(x)\rvert^{2}\exp(-U(x)/\varepsilon)\,\mathrm{d}x$
subject to $\tilde{q}=0$ on $\partial A$ and $\tilde{q}=1$ on $\partial B$ [3]. It is not clear how the value of the functional $J(\tilde{q})$ relates to the discrepancy between the distributions of exact and approximate reactive trajectories. In that sense, it is not a such a natural loss function as our relative entropy. In Section 4.2, we devise a model assessment procedure that compares approximate committor functions based on relative entropy differences as in the Bayes and Akaike information criteria.
@@
All applications of our results demand that extraordinary care be taken to correctly handle singularities. For example, since $q=0$ on $\partial A$, the drift $2\varepsilon\nabla\log q(Y_{t})$ in (4) is singular on $\partial A$, and standard numerical integrators may fail to converge. We have devised a weakly convergent operator splitting scheme for integrating (2) in the case where $\partial A$ is planar and the approximate committor function meets a certain boundary condition on $\partial A$. We will publish a numerical analysis of this integrator in a companion to the present article. We have not yet been able to devise a practical, provably convergent integrator that correctly handles the singularity on $\partial A$ in all cases. We also note that the change of measure formula (11) involves an integral that may be singular, but we have devised a general means of treating the singularity and correctly computing the integral; cf. Section 5.
2. A Digest of Transition Path Theory
@@
In this section, we present a brief review of transition path theory [5, 13]. As above, let $U:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a potential energy, describing for example a molecular system. Let $\varepsilon=k_{B}T>0$ be a temperature parameter, and suppose that $X_{t}$ evolves according to the overdamped Langevin dynamics
$\mathrm{d}X_{t}=-\nabla U(X_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}.$ (3)
@@
In transition path theory, to characterize rare transitions of the overdamped Langevin dynamics, one first chooses a disjoint pair of subsets $A$ and $B$ of $\mathbb{R}^{d}$. In applications, $A$ and $B$ would usually be associated with metastable states, e.g. they might be sets modeling different conformations of a biomolecule. In addition to the sets $A$ and $B$, we find it convenient to define the transition region
@@
Figure 1. Schematic illustration of reactive trajectories, entrance times, and exit times. The black line depicts a trajectory of overdamped Langevin dynamics. The bold red segments depict reactive trajectories. Here, $\tau_{B,0}$ is the first hitting time of $B$, $\tau_{A,1}$ is the first hitting time of $A$ after $\tau_{B,0}$, $\tau_{B,1}$ is the first hitting time of $B$ after $\tau_{A,1}$, and so forth. The $k$’th exit time $\sigma_{A,k}$ is the last time before $\tau_{B,k}$ that the process was in $A$.
Figure 2. A trajectory of the overdamped Langevin dynamics for a simple two-dimensional model potential. The blue curves are contours of the potential. The trajectory depicted here was initialized from the Boltzmann distribution, i.e. in equilibrium. Its starting point happened to be in the set $A$ on the left side of the figure. The trajectory was terminated on hitting the boundary of $B$ for the first time. The bold red curve depicts the end of the trajectory from the last time it left $A$ to the first time it entered $B$; i.e. the reactive part of the trajectory. The faint red curve depicts the beginning of the trajectory up to the last time it left $A$.
@@
$\mathrm{d}Y_{t}=-\nabla U(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla\log q(Y_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}$ (4)
@@
$Y_{0}\sim\frac{1}{\nu}\lvert\nabla q(x)\rvert\exp(-U(x)/\varepsilon)\,\mathrm{d}S(x).$ (5)
@@
$\nu=\int_{\partial A}\lvert\nabla q(x)\rvert\exp(-U(x)/\varepsilon)\,\mathrm{d}S(x).$
Observe that the transition path equation is the overdamped Langevin dynamics with the additional drift term $2\varepsilon\nabla\log q(X_{t})\,\mathrm{d}t$. On $\partial A$, we have $q=0$, and by Lemma 1 below, $\lvert\nabla q\rvert>0$. Therefore, the drift $\nabla\log q=\frac{\nabla q}{q}$ is singular on $\partial A$. Despite the singularity, Lu and Nolen showed that a unique strong solution of (4) exists for any initial distribution, even distributions supported on $\partial A$. Moreover, with probability one, $X_{t}\notin A$ for $t>0$. In effect, the singular drift repels trajectories from $A$, forcing transitions to occur.
Remark 1.
@@
$\begin{cases}Lq=0&\text{ in }\mathscr{T},\\ q=0&\text{ on }\partial A,\\ q=1&\text{ on }\partial B,\end{cases}$ (6)
@@
is the generator of the overdamped Langevin dynamics [5]. Therefore, under our smoothness assumptions on $U$ and $\mathscr{T}$, elliptic regularity, the Hopf lemma, and the maximum principle imply the properties of the committor listed in Lemma 1 below.
Assumption 3.
@@
Lemma 1.
Under Assumptions 2 and 3, the forward committor function $q:\mathscr{T}\rightarrow[0,1]$ extends to an infinitely differentiable function defined on an open set containing $\partial A$ and $\partial B$. By abuse of notation, we let $q$ refer to both the committor function and the extension. For all $x\in\mathring{\mathscr{T}}$,
@@
Proof.
By elliptic regularity, there exists an infinitely differentiable extension of $q$ to an open set containing $\overline{\mathscr{T}}$. We have $\nabla q(x)\cdot n(x)>0$ for $x\in\partial A$ by the Hopf Lemma, and $q(x)>0$ in the interior of $\mathscr{T}$ by the maximum principle. Moreover, since $q$ is constant on $\partial A$ and $\nabla q(x)\cdot n(x)>0$, $\nabla q(x)=\lvert\nabla q(x)\rvert n(x)$. ∎
Remark 2.
It is conventional to take the domain of the committor to be $\mathbb{R}^{d}$ instead of $\mathscr{T}$, letting $q=0$ on $A$ and $q=1$ on $B$. This extension of $q$ does not coincide with the smooth extension described in Lemma 1. Any smooth extension must take negative values at some points in the interior of $A$ since $\nabla q\cdot n>0$ and $q=0$ on $\partial A$. Moreover, observe that for our smooth extension, $Lq=0$ everywhere in $\bar{\mathscr{T}}$, including on $\partial A$. In the original work on transition state theory, $Lq$ was instead understood as a distribution supported on $\partial A\cup\partial B$. We require a smooth extension only to simplify the notation in certain proofs. None of our results depends on a particular choice of extension.
3. A Change of Measure Formula for Approximations to the Transition Path Process
We analyze the errors that result when one substitutes an approximation $\tilde{q}$ of the committor in the transition path equation (4). To be precise, in Theorem 1 below, we present a change of measure formula relating the distribution of solutions of the approximate transition path equation
$\mathrm{d}Y_{t}=-\nabla U(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla\log\tilde{q}(Y_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}$ (7)
with the exact transition path process. To derive the change of measure, motivated by Lemma 1, we impose the following assumptions on the approximate committor $\tilde{q}$.
Assumption 4.

### v2

# Relative Entropy Methods for the Approximation of Reactive Trajectories thanks: BvK was supported by NSF DMS-2012207

@@
Abstract
Motivated by challenges arising in molecular simulation, we study reactive trajectories of the overdamped Langevin dynamics, i.e. trajectories observed as they pass from a set $A$ corresponding to the reagents of a chemical reaction to a set $B$ corresponding to the products. Reactive trajectories are known to have the same distribution as trajectories of the overdamped Langevin dynamics biased by a singular drift related to the committor function. In this work, we assess the effect of replacing the exact singular drift with an approximation based on an approximate committor function. We derive a convenient formula for the relative entropy between the distributions of exact and approximate reactive trajectories, and we propose a stochastic gradient descent method for minimizing the entropy to train an approximate committor function on the fly while computing reactive trajectories. We also devise a model assessment procedure for comparing the qualities of different approximations to the committor function based on the relative entropy.
1 Introduction
For many molecular systems, the most interesting phenomena relate to extremely rare transitions between long-lived states, e.g. transitions between folded states of a protein. Such systems are called metastable. It may be infeasible to perform a simulation long enough to observe even one rare transition in a metastable system. Therefore, a wide variety of methods have been devised to compute statistics of rare transitions by means other than direct simulation. Some of these methods, like adaptive multilevel splitting [9, 10], forward flux sampling [1], and transition path sampling [6] produce samples of transition paths or fragments of transition paths. Others produce a representative path, for example the minimizer of the Wentzell–Friedlin action, which is roughly speaking the mode of the distribution of transition paths in the limit of low temperature [29]; see [13] for a survey. Finally, one can compute many important statistics of transition paths in terms of solutions of partial differential equations using transition path theory [12, 13]. By various means, these methods efficiently explore regions associated with rare transitions.
Motivated by the challenge of simulating metastable systems, we study reactive trajectories of the overdamped Langevin dynamics
(1) $\mathrm{d}X_{t}=-\nabla U(X_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t},$
which are trajectories observed as they pass from a set $A$ corresponding to the reagents of a chemical reaction to a set $B$ corresponding to the products; cf. Figures 1 and 2. In [24], building on the transition path theory of E and Vanden-Eijnden [12], Lu and Nolen showed that reactive trajectories have the same distribution as the unique strong solution of the transition path equation
(2) $\mathrm{d}Y_{t}=-\nabla U(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla\log q(Y_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}$
with initial condition
(3) $Y_{0}\sim\frac{1}{\zeta}\lvert\nabla q(x)\rvert\exp(-U(x)/\varepsilon)\,\mathrm{d}S_{A}(x).$
Here,
$q(x)=\mathds{P}[\tau_{B}<\tau_{A}|Y_{0}=x]$
@@ 1 Introduction
is the committor function, i.e. the probability that a trajectory of the overdamped Langevin dynamics (1) with $Y_{0}=x$ will hit $B$ before $A$, $S_{A}$ is the surface measure on $\partial A$, and $\zeta$ is a normalizing constant. The drift term $2\varepsilon\nabla\log q(X_{t})$ in (2) is singular on $\partial A$, since $q$ is zero on $\partial A$, and it forces trajectories away from $A$ and towards $B$ so that the lengths of reactive trajectories will typically be much shorter than the intervals of time between reactions; cf. Figure 2. That is, the singular drift is a bias that forces rare transitions to occur quickly and according to the correct distribution.
[…]
To explain why we are interested in methods of this type, we wish to make two comparisons with other classes of methods in the computational chemistry literature. First, adaptive multilevel splitting [9] computes a sample from the exact distribution of reactive trajectories initiated from a fixed starting point near the boundary of $A$ by branching and killing trajectory fragments. Our methods produce approximate reactive trajectories, but since we calculate the committor and trajectories together, we can ensure that the initial points have at least approximately the reactive flux distribution (3). [^2] We note that, in [23], Lopes and Lelièvre propose a very promising means of determining the right distribution of initial points for adaptive multilevel splitting, but their method requires the choice of a surface close to $A$ having certain properties that might be difficult to verify in practice. Our approach does not require the identification of any such surface. Moreover, unlike adaptive multilevel splitting, our approach does not require the explicit specification of a reaction coordinate. In fact, while computing reactive trajectories, we solve for the committor, which is in some respects the ideal reaction coordinate [10].
[…]
To summarize, our proposed relative entropy optimization method computes both reactive trajectories and the committor, with error estimates, with at least approximately the correct distribution of initial points, and without requiring an explicit choice of collocation points or a reaction coordinate. Instead, the user must specify an initial approximation $\tilde{q}_{0}$ to the committor function. We discuss the choice of $\tilde{q}_{0}$ in Section 6. Here, we note only that one could choose $\tilde{q}_{0}$ so that the transition path process with $\tilde{q}_{0}$ in place of $q$ imposes a biasing force pushing on a small number of atoms to induce transitions to occur as in steered molecular dynamics [17].
In addition to optimizing the relative entropy, we demonstrate two other potential applications of our theoretical results: The assessment of models of the committor and importance sampling. For high-dimensional systems, the committor and related quantities such as the density and current of transition paths may be difficult to interpret without coarse-graining, i.e. dimensionality reduction. We suggest that one could select or optimize coarse models of the committor based on the relative entropy; cf. Section 6. We also note that one can use importance sampling to correct for errors in the distribution of reactive trajectories generated based on an approximate committor. However, the change of measure involves an exponential average, and experience with similar importance sampling methods based on Girsanov’s theorem suggests that such an approach would not be reliable for poor approximations of the committor; cf. Section 4.4. Nonetheless, importance sampling yields good results for our simple test problem in Section 6.5.
@@
We warn the reader that all theoretical and computational applications of our results demand that care be taken to correctly handle singularities. For example, since the drift is singular standard numerical integrators may fail. We have proposed an operator splitting scheme for integrating (2) in the case where $\partial A$ is planar. We do not prove convergence of the integrator, nor do we propose an integrator that works when $\partial A$ is curved, although we do grant that one would often want a curved boundary. We also note that the change of measure formula (12) involves an integral whose integrand may be singular; cf. Section 5.2.
Finally, we note that we are only able to prove the existence of the gradient $\nabla_{\theta}D_{\mathrm{KL}}(\mathds{P}_{q_{\theta}}\|\mathds{Q})$ under rather stringent conditions; cf. Assumption 4.3. We suspect, but cannot prove, that the relative entropy is differentiable under weaker conditions. In any case, our method of training the committor by stochastic gradient descent appears to work well even when Assumption 4.3 fails; cf. Section 6.4. Note that we require Assumption 4.3 only to prove differentiability, not to derive the change of measure formula in Theorem 3.11 or to prove the weak law of large numbers for the estimator of relative entropy differences in Section 4.2.
Outline. In Section 2, we briefly review transition path theory. In Section 3, we derive the change of measure between the exact distribution of transition paths and the distribution of (2) with an approximate committor $\tilde{q}$ in place of $q$. In Section 4, we explain how to estimate relative entropy differences and gradients, and we consider the possibility of importance sampling. In Section 5, we address some practical obstacles to the implementation of our methods, including strategies for the representation of approximate committor functions that satisfy Assumption 3 and Assumption 4.3. Finally, in Section 6, we apply our methods to a two-dimensional toy model. First, we propose an integrator capable of handling the singular drift. We then assess the quality of a coarse approximation of a committor, train an improved approximation by stochastic gradient descent, demonstrate the use of importance sampling to correct for errors in the committor, and study the convergence of the integrator.
2 A Digest of Transition Path Theory
@@
In this section, we present a brief review of transition path theory [12, 24]. Let $U:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a potential energy, describing for example a molecular system. Let $\varepsilon=k_{B}T>0$ be a temperature parameter, and suppose that $X_{t}$ evolves according to the overdamped Langevin dynamics
(4) $\mathrm{d}X_{t}=-\nabla U(X_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}.$
@@
In transition path theory, to characterize rare transitions of the overdamped Langevin dynamics, one first chooses a disjoint pair of subsets $A$ and $B$ of $\mathbb{R}^{d}$. In applications, $A$ and $B$ would usually be associated with metastable states, e.g. they might be sets modeling different conformations of a biomolecule or the reagents and products of a chemical reaction. In addition to the sets $A$ and $B$, we find it convenient to define the transition region
@@
Figure 1: Schematic illustration of reactive trajectories, entrance times, and exit times. The orange line depicts a trajectory of the overdamped Langevin dynamics. The bold red segments depict reactive trajectories. Here, $\tau_{B,0}$ is the first hitting time of $B$, $\tau_{A,1}$ is the first hitting time of AA after $\tau_{B,0}$, $\tau_{B,1}$ is the first hitting time of $B$ after $\tau_{A,1}$, and so forth. The $k$’th exit time $\sigma_{A,k}$ is the last time before $\tau_{B,k}$ that the process was in $A$.
Figure 2: Reactive trajectories. On the left, we have a single long trajectory of the overdamped Langevin dynamics for the simple two-dimensional model potential introduced in Section 6. The blue curves are contours of the potential. The trajectory was initialized from the Boltzmann distribution. Its starting point happened to be in the set AA on the left side of the figure. The trajectory was terminated on hitting the boundary of $B$ for the first time. The bold red curve depicts the end of the trajectory from the last time it left AA to the first time it entered $B$, i.e. the reactive part of the trajectory. The faint orange curve depicts the beginning of the trajectory up to the last time it left AA. On the right, we have three approximate reactive trajectories computed by integrating the transition path process for the approximate committor $q_{\rm{sgd}}$ calculated by gradient descent in Section 6.4. The blue curves are contours of the effective potential $U-2\varepsilon\log q_{\rm{sgd}}$ for the transition path process, which is singular on $\partial A$. The red and pink curves are the reactive trajectories.
@@
(5) $\mathrm{d}Y_{t}=-\nabla U(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla\log q(Y_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}$
@@
(6) $Y_{0}\sim\frac{1}{\zeta}\lvert\nabla q(x)\rvert\exp(-U(x)/\varepsilon)\,\mathrm{d}S(x).$
@@
$\zeta=\int_{\partial A}\lvert\nabla q(x)\rvert\exp(-U(x)/\varepsilon)\,\mathrm{d}S(x).$
Observe that the transition path equation is the overdamped Langevin dynamics with the additional drift term $2\varepsilon\nabla\log q(X_{t})\,\mathrm{d}t$. On $\partial A$, we have $q=0$, and by Lemma 2.2 below, $\lvert\nabla q\rvert>0$. Therefore, the drift $\nabla\log q=\frac{\nabla q}{q}$ is singular on $\partial A$. Despite the singularity, Lu and Nolen showed that a unique strong solution of (5) exists for any initial distribution, even distributions supported on $\partial A$. Moreover, with probability one, $X_{t}\notin A$ for $t>0$. In effect, the singular drift repels trajectories from $A$, forcing transitions to occur.
Remark 2.1.
@@
(7) $\begin{cases}Lq=0&\text{ in }\mathscr{T},\\ q=0&\text{ on }\partial A,\\ q=1&\text{ on }\partial B,\end{cases}$
@@
is the generator of the overdamped Langevin dynamics [12]. Therefore, under our smoothness assumptions on $U$ and $\mathscr{T}$, elliptic regularity, the Hopf lemma, and the maximum principle imply the properties of the committor listed in Lemma 2.2 below.
{assumption}
@@
Lemma 2.2.
Under Assumptions 2 and 2, the forward committor function $q:\mathscr{T}\rightarrow[0,1]$ extends to an infinitely differentiable function defined on an open set containing $\partial A$ and $\partial B$. By abuse of notation, we let $q$ refer to both the committor function and the extension. For all $x\in\mathring{\mathscr{T}}$,
@@
Proof 2.3.
By elliptic regularity, there exists an infinitely differentiable extension of $q$ to an open set containing $\overline{\mathscr{T}}$. We have $\nabla q(x)\cdot n(x)>0$ for $x\in\partial A$ by the Hopf Lemma, and $q(x)>0$ in the interior of $\mathscr{T}$ by the maximum principle. Moreover, since $q$ is constant on $\partial A$ and $\nabla q(x)\cdot n(x)>0$, $\nabla q(x)=\lvert\nabla q(x)\rvert n(x)$.
Remark 2.4.
It is conventional to take the domain of the committor to be $\mathbb{R}^{d}$ instead of $\mathscr{T}$, letting $q=0$ on $A$ and $q=1$ on $B$. This extension of $q$ does not coincide with the smooth extension described in Lemma 2.2. Any smooth extension must take negative values at some points in the interior of $A$ since $\nabla q\cdot n>0$ and $q=0$ on $\partial A$. Moreover, observe that for our smooth extension, $Lq=0$ everywhere in $\bar{\mathscr{T}}$, including on $\partial A$. In the original work on transition state theory, $Lq$ was instead understood as a distribution supported on $\partial A\cup\partial B$. We require a smooth extension only to simplify the notation in certain proofs. None of our results depends on a particular choice of extension.
3 A Change of Measure Formula for Approximations to the Transition Path Process
We analyze the errors that result when one substitutes an approximation $\tilde{q}$ of the committor in the transition path equation (5). To be precise, in Theorem 3.11 below, we present a change of measure formula relating the distribution of solutions of the approximate transition path equation
(8) $\mathrm{d}Y_{t}=-\nabla U(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla\log\tilde{q}(Y_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}$
with the exact transition path process. To derive the change of measure, motivated by Lemma 2.2, we impose the following assumptions on the approximate committor $\tilde{q}$.
{assumption}